# IGNITE: Individualized GeNeration of Imputations in Time-series Electronic health records

### v1

@@ 1.2 Reconstruction Task
Other than measuring the performance of IGNITE in predictive modelling tasks, we also evaluate its ability to recover and reconstruct randomly introduced missing values (i.e., missing at completely random). In the reconstruction experiments, IGNITE consistently outperformed the other baselines across the two metrics, Root-Mean-Square Error (RMSE) and Mean Absolute Error (MAE) for each patient. The metrics shown in Table 12, are reported in the mean and standard error of the errors across all the patients in the test set. Specifically, IGNITE had a RMSE of 0.100 (0.038), followed by MICE with 0.109 (0.059) for the setting where 10% missingness was introduced in the PhysioNet 2012 data. Similar trends were observed in the 20% and 50% experiments, where IGNITE achieved the lowest reconstruction error and LOCF achieved the highest error. It is worth noting that the error in IGNITE generally remained stable despite increasing the percentage of missingness, indicating its robustness in reconstructing data with high missingness. We showcase the full results for the reconstruction experiments on HiRID and eICU datasets in the Supplementary materials E, where IGNITE consistently outperformed all the baselines. Sample visualization of the reconstructed values via top-performing models is shown in Figure 2. In Figure 2a and Figure 2b, we show the imputations generated by the various models for two patients with different types of missingness ie: with feature-wise and sample-wise missingenss, where IGNITE was the model with the lower error and best at recovering masked values. In Figure 2c, we see that IGNITE and the other deep learning models generated realistic values. However, LOCF imputation fails as there are no observed values. The other methods’ reconstructions are shown in Supplementary materials F .
Table 5: Performance of the various baselines of the reconstruction task on the PhysioNet 2012 dataset. The results are calculated across all the patients in the test set and reported in terms of RMSE and MAE with mean and standard deviation.
@@
LOCF 0.183 (0.065) 0.192 (0.051) 0.202 (0.036) 0.106 (0.040) 0.108 (0.034) 0.115 (0.028)
MICE 0.109 (0.046) 0.112 (0.037) 0.122 (0.034) 0.066 (0.023) 0.066 (0.019) 0.066 (0.017)
GP-VAE 0.395 (0.067) 0.398 (0.521) 0.398 (0.044) 0.317 (0.056) 0.318 (0.045) 0.317 (0.038)
Transformer 0.119 (0.059) 0.128 (0.05) 0.123 (0.039) 0.068 (0.028) 0.07 (0.025) 0.068 (0.021)
BRITS 0.115 (0.054) 0.116 (0.044) 0.118 (0.038) 0.066 (0.027) 0.066 (0.022) 0.065 (0.02)
SAITS 0.122 (0.058) 0.127 (0.051) 0.123 (0.038) 0.071 (0.029) 0.069 (0.025) 0.067 (0.021)
@@
Figure 2: Visualized imputations for patients from PhysioNet 2012 dataset. In a and b, we show examples of patients with different types of missingness. The original observed values are shown as black. We further masked 50% of the observed values and considered they are ground truth as shown in green. Various imputation methods are compared with respect to the ground truth.
@@ 1.3 Ablation Study
To demonstrate the added value of each of the model’s components, we conducted an ablation study on the Physionet 2012 challenge dataset. The studied components were mainly the time-series conditional component as well as the IMM mask used in IGNITE. As presented in Table 6, we observe consistent performance gains in terms of AUROC as new components are added. The highest performance gain was attributed to adding IMM to the model, where the final model achieved 0.835 of AUROC.
Table 6: Performance when ablating various components of the proposed models, reported on the PhysioNet 2012 dataset in terms of AUROC
Ablations AUROC
IGNITE (Dual VAE, No Condition) 0.818
IGNITE (Dual VAE, Condition) 0.823
IGNITE (Dual VAE, Condition, IMM) 0.830
IGNITE (Dual VAE, Condition, IMM, MIT loss) 0.832
IGNITE (Dual VAE, Condition, IMM, MIT loss, discriminator) 0.835
@@ Discussion
In this work, we proposed a deep end-to-end generative model that synthesizes personalized data in highly sparse and irregularly sampled or even completely missing EHR conditioned on treatments and demographic characteristics, as well as the individual level of missingness patterns. We note that our model, IGNITE, outperformed the baseline models on the full population downstream mortality prediction task in terms of AUROC and AUPRC. Unlike previous works, we proposed a new framework for the evaluation of imputation models trained on various types of missingness across features and samples in time-series data. In the proposed downstream task evaluation, IGNITE showed consistent robustness despite varying the percentage of feature-wise and sample-wise missingness. By augmenting our training with a novel individualized missingness mask, we believe IGNITE captured better dynamics and underlying indicators of the patient’s health status.
@@
This paper has several strengths. First, this is the first work to introduce a novel missingness mask that represents the measurement patterns and frequencies in an individualized way. The proposed IMM mask showed high utility in imputation tasks, we believe such masks can be potentially utilized in other diverse time-series tasks such as prediction Han et al., (2019), detection Blázquez-García et al., (2021), clustering Aghabozorgi et al., (2015) and more. Second, the proposed work is the first to investigate feature-wise missingness and sample-wise missingness in observed healthcare records, bringing a new understanding of missingness patterns in real-world data. While several works investigated theoretical missingness assumptions such as MCAR, MAR and MNAR Curran et al., (1998), such missingness mechanisms are hard to prove and remain far from patterns observed in real-world health data. We found a correlation between the positive outcome prevalence and sampling frequency by investigating the percentage of outcome prevalence across populations with features that were never observed and those with different overall sample missingness. There was a significantly higher percentage of positive outcomes in patients with $\leq$25of features never observed across all three datasets, where the outcome prevalence more than doubled compared to those with higher feature-wise missingness. Similar trends were also observed in sample-wise missingness, which suggests that the missingness pattern is personalized and could inform the underlying patient health status for various personalized medical applications. Imputation methods that consider individual missingness patterns, such as IGNITE, could lead to more accurate predictions and tailored treatments. Moreover, our feature-wise missingness experiments showcased IGNITE’s power to generate features that were never observed in patients. In traditional statistical approaches, many patients with such never-observed features are generally omitted/removed leading to a waste of data or a reduction in model performance. However, with IGNITE, such features can be generated based on the other readily available information of an individual. We believe that a flexible generative approach proposed by IGNITE can open doors for a new way of rethinking missingness imputation as a generative task, making various new applications utilizing generative models.
@@
Another strength of our proposed work is that IGNITE is designed to generate imputations conditioned on individualized patient observed and missingness patterns and dynamics, making the generated imputations of high quality and more aligned with personalized medicine applications. Generating digital twins to drive personalized medicine applications requires understanding and representing the patient data beyond the observed values, which IGNITE incorporates in its generative process. Lastly, IGNITE was trained and tested across three large-scale EHRs, outperforming state-of-the-art imputation models across a series of experiments. Other experiments, such as the reconstruction experiments, demonstrate IGNITE’s robustness and ability to recover original patient values even when random missingness is introduced to patient records. This robust performance across datasets and experimental settings showcases IGNITE’s robustness in learning individualized representations and generating high-quality imputations.
Overall, this study has some limitations. IGNITE and the other imputation models were only tested on retrospective ICU datasets. In future works, we plan to investigate IGNITE’s ability to learn individualized patient representations with higher sparsity and missingness patterns in primary care and wearable data. Furthermore, while our work expands on the evaluation metrics proposed in the literature, IGNITE was only evaluated in prediction tasks involving mortality as limited by the available datasets such as Physionet 2012. In future works, we aim to investigate the impact of imputations on various tasks in healthcare settings, such as treatment recommendation and phenotyping. Another area of future work involves the theoretical analysis of the various missingness patterns in the data to better understand the strengths and limitations of the plethora of time-series techniques.
@@ 1.4 Notations
We denote a multivariate time-series EHRs dataset as $\mathbf{D}=\left\{\left(\mathbf{x}_{i,1:T_{i}}\right)\right\}_{i=1}^{N}$, which includes a set of individual patient records indexed by $i\in\{1,2,\ldots N\}$. Each patient record comprises two main components: time-series and static components. The time-series component includes continuous time-series of physiological data and the discrete time-series of treatment data, recorded over time steps $\mathbf{T}=(\mathbf{t}_{1},\ldots,{t}_{N})$. Each record has a corresponding static outcome label $\mathbf{y}=(\mathbf{y}_{1},\ldots,{y}_{N})$ and static demographic information, indicating the age and sex of the patient. For each individual patient time-series matrix, we extract a binary mask indicating the missingness of each value, where 0 and 1 denote the missing value and observed value, respectively.
@@ 1.5.1 Dual Variational-AutoEncoders with Dual-Stage Attention
We utilize a pair of VAEs, each with an encoder and a decoder, to map the multi-variate time-series into reversible low-dimensional dense representations with no missingness. We use long short-term memory (LSTM) Hochreiter and Schmidhuber, (1997) neural networks for the architecture of the VAE. The first VAE is trained to reconstruct the multivariate time-series by calculating the evidence lower bound (ELBO) on the observed values only and imputing the missing values with zero, an approach adopted by (Nazabal et al.,, 2020; Fortuin et al.,, 2020). The second VAE, on the other hand, generates the data based on the ELBO calculated on the full data that is first imputed with the last value carried forward and masked with the Individualized Missingness Mask (IMM) that is discussed in section 1.5.2. We refer to the first and second VAE as $VAE^{\mathcal{OO}}$and $VAE^{\mathcal{IMM}}$, respectively. Each LSTM encoder is complemented with input attention highlighting feature importance. Specifically, the attention component computes the weights for each feature conditioned on the encoder’s hidden state of the previous time step. The computed weights are obtained using a deterministic fully connected network, by referring to the previous hidden state $\mathbf{h}_{t-1}$ and the cell state $\mathbf{s}_{t-1}$ in the encoder LSTM unit with:
@@
where $\alpha_{t}^{k}$ is the attention weight of feature $k$ at time $t$. Learnable parameters are denoted in the form of $\mathbf{v}_{e}\in\mathbb{R}^{T},\mathbf{W}_{e}\in\mathbb{R}^{T\times 2m}$ and $\mathbf{U}_{e}\in\mathbb{R}^{T\times T}$. The learnt attention weights are passed to a softmax function, followed by a dense network trained with the encoder parameters. To this end, the learned weights are multiplied by the input features as shown in
@@
The LSTM encoder unit are fed with $\tilde{\mathbf{x}}_{t}$ instead of ${\mathbf{x}}_{t}$, giving more attention to important features to encode a better representation of the time-series input. After encoding the data with the feature-wise attention, a temporal attention mechanism is used in the decoder to select relevant encoder hidden states across all time steps. To do so, an attention weight is computed for each encoder hidden state at time $t$. The calculation of the attention weight is based on the previous hidden state at time $t-1$. Specifically, the attention weight of each encoder hidden state at time $t$ is calculated based on the previous decoder hidden state $\mathbf{d}_{t-1}\in\mathbb{R}^{p}$ and the cell state of the LSTM unit $\mathbf{s}_{t-1}^{\prime}\in\mathbb{R}^{p}$ with
@@
where $\left[\mathbf{d}_{t-1};\mathbf{s}_{t-1}^{\prime}\right]\in\mathbb{R}^{2p}$ is a concatenation of the previous hidden state and cell state of the LSTM unit. $\mathbf{v}_{d}\in\mathbb{R}^{m}$, $\mathbf{W}_{d}\in\mathbb{R}^{m\times 2p}$ and $\mathbf{U}_{d}\in\mathbb{R}^{m\times m}$ are parameters to learn. The attention weight $\beta_{t}^{i}$ represents the importance of the $i$-th encoder hidden state at time $t$. Since each encoder hidden state $\mathbf{h}_{i}$ is mapped to a temporal component of the input, the attention mechanism computes the context vector $\mathbf{c}_{t}$ as a weighted sum of all the encoder hidden states $\left\{\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{T}\right\}$,
@@
We note that the context vector $\mathbf{c}_{t}$ is distinct at each time step. Once we get the weighted summed context vectors. The decoded data is based on the stacked weighted content vectors, allowing for better representation over time.

### v2

@@ 1.2 Reconstruction Task
Other than measuring the performance of IGNITE in predictive modelling tasks, we also evaluate its ability to recover and reconstruct randomly introduced missing values (i.e., missing at completely random). In the reconstruction experiments, IGNITE consistently outperformed the other baselines across the two metrics, Root-Mean-Square Error (RMSE) and Mean Absolute Error (MAE) for each patient. The metrics shown in Table 13, are reported in the mean and standard error of the errors across all the patients in the test set. Specifically, IGNITE had an RMSE of 0.100 (0.038), followed by MICE with 0.109 (0.059) for the setting where 10% missingness was introduced in the PhysioNet 2012 data. Similar trends were observed in the 20% and 50% experiments, where IGNITE achieved the lowest reconstruction error and LOCF achieved the highest error. It is worth noting that the error in IGNITE generally remained stable despite increasing the percentage of missingness, indicating its robustness in reconstructing data with high missingness. We showcase the full results for the reconstruction experiments on HiRID and eICU datasets in the Supplementary materials F, where IGNITE consistently outperformed all the baselines. In addition to its methodological contributions and high performance in reconstruction tasks, IGNITE significantly boosts prediction accuracy by improving how imputed data align with actual patient statuses — a crucial factor in effective clinical decision-making. By accurately imputing missing data in a way that reflects real-time changes in patient health, IGNITE preserves essential temporal trends and patient-specific dynamics within the EHR data. We show example patient imputations with different patient states in sample visualization of the reconstructed values via top-performing models shown in Figure 2. In Figure 2a and Figure 2b, we show the imputations generated by the various models for two patients with different types of missingness ie: with feature-wise and sample-wise missingenss, where IGNITE was the model with the lower error and best at recovering masked values. In Figure 2b, we see that IGNITE and the other deep learning models generated realistic values. However, LOCF imputation fails as there are no observed values. The other methods’ reconstructions are shown in Supplementary materials G. Furthermore, In Figure3, We present a comparison of imputed values using IGNITE against BRITS and MICE both with strong performance. In this section, we showcase imputed values for three key variables, with additional results provided in the Supplementary Material H. Our results demonstrate that IGNITE is the only model that imputes urine output within the true range and trend, as seen in a case where the patient’s urine output decreases prior to death. In contrast, BRITS and MICE show an increasing trend, far from the ground truth. A similar observation is made for HCO3, where IGNITE is the only model to predict values close to the ground truth. For FiO2, typically flat around 0.5 for patients on mechanical ventilation, IGNITE accurately captures this, providing realistic imputations. These examples highlight that IGNITE not only improves predictive performance but also generates clinically realistic imputations.
Table 5: Performance of the various baselines of the reconstruction task on the PhysioNet 2012 dataset. The results are calculated across all the patients in the test set and reported in terms of RMSE and MAE with mean and standard deviation. $\ast$ indicates that IGNITE is statistically significant ($p<0.05$) when compared to the corresponding benchmark.
@@
LOCF 0.183 (0.065)* 0.192 (0.051)* 0.202 (0.036)* 0.106 (0.040)* 0.108 (0.034)* 0.115 (0.028)*
MICE 0.109 (0.046)* 0.112 (0.037) 0.122 (0.034) * 0.066 (0.023) 0.066 (0.019) 0.066 (0.017)
GP-VAE 0.395 (0.067)* 0.398 (0.521)* 0.398 (0.044)* 0.317 (0.056) * 0.318 (0.045)* 0.317 (0.038)*
Transformer 0.119 (0.059)* 0.128 (0.05)* 0.123 (0.039)* 0.068 (0.028) 0.07 (0.025)* 0.068 (0.021)*
BRITS 0.115 (0.054)* 0.116 (0.044)* 0.118 (0.038)* 0.066 (0.027) 0.066 (0.022) 0.065 (0.02)
TimesNet 0.266 (0.158)* 0.263 (0.045)* 0.266 (0.144)* 0.071 (0.026)* 0.069 (0.021)* 0.079* (0.027 )
SAITS 0.122 (0.058)* 0.127 (0.051)* 0.123 (0.038)* 0.071 (0.029)* 0.069 (0.025)* 0.067 (0.021)
@@
Figure 2: Visualized imputations for patients from PhysioNet 2012 dataset. In a and b, we show examples of patients with different types of missingness. The original observed values are shown as black. In a, we show an example of a patient with sample-wise missingness, while in b, we show a patient with feature-wise missingness indicating no observed measurements for that feature across all time-steps. We further masked 50% of the observed values and considered them to be the ground truth as shown in green. Various imputation methods are compared with respect to ground truth.
(a) Urine Output
(b) FiO2
(c) HCO3
Figure 3: Visualized imputations for a dead patient, where we compare IGNITE to two models, namely BRITS and MICE. The original observed values are shown as black and the masked values that are considered ground truth are shown as green. We present the rest of the imputed variables for this patient in the Supplementary Information H.
@@ 1.3 Ablation Study
To demonstrate the added value of each of the model’s components, we conducted an ablation study on the three datasets The studied components were mainly the time-series conditional component as well as the IMM mask used in IGNITE. As presented in Table 6, we observe consistent performance gains in terms of AUROC as new components are added across the three datasets. The highest performance gain was attributed to adding IMM to the model, where the final model achieved 0.835, 0.774 and 0,968 of AUROC for PhysioNet 2012, eICU and HiRID respectively.
Table 6: Performance when ablating various components of the proposed models, reported in terms of AUROC
Ablations PhysioNet 2012 eICU HiRID
IGNITE (Dual VAE, No Condition) 0.818 0.746 0.927
IGNITE (Dual VAE, Condition) 0.823 0.754 0.933
IGNITE (Dual VAE, Condition, IMM) 0.830 0.762 0.952
IGNITE (Dual VAE, Condition, IMM, MIT loss) 0.832 0.765 0.965
IGNITE (Dual VAE, Condition, IMM, MIT loss, discriminator) 0.835 0.774 0.968
@@ Discussion
In this work, we proposed a deep end-to-end generative model that synthesizes personalized data in highly sparse and irregularly sampled or even completely missing EHR conditioned on treatments and demographic characteristics, as well as the individual level of missingness patterns. We note that our model, IGNITE, outperformed the baseline models in the full population downstream mortality prediction task in terms of AUROC and AUPRC. Unlike previous works, we proposed a new framework for the evaluation of imputation models trained on various types of missingness across features and samples in time-series data. In the proposed downstream task evaluation, IGNITE showed consistent robustness despite varying the percentage of feature-wise and sample-wise missingness. By augmenting our training with a novel individualized missingness mask, we believe that IGNITE captured better dynamics and underlying indicators of the patient’s health status. Unlike traditional imputation methods that often assume a specific missingness mechanism, IGNITE is designed to be adaptable to various types of missing data scenarios commonly encountered in healthcare datasets. This adaptability improves the utility of the model in a wide range of clinical applications beyond the datasets used in this study.
@@
This paper has several strengths. First, this is the first work to introduce a novel missingness mask that represents measurement patterns and frequencies in an individualized way. The proposed IMM mask showed high utility in imputation tasks, we believe such masks can be potentially utilized in other diverse time-series tasks such as prediction Han et al., (2019), detection Blázquez-García et al., (2021), clustering Aghabozorgi et al., (2015) and more. Second, the proposed work is the first to investigate feature-wise missingness and sample-wise missingness in observed healthcare records, bringing a new understanding of missingness patterns in real-world data. Although several works investigated theoretical missingness assumptions such as MCAR, MAR and MNAR Curran et al., (1998), such missingness mechanisms are hard to prove and remain far from patterns observed in real-world health data. We found a correlation between the prevalence of positive outcomes and sampling frequency by investigating the percentage of outcome prevalence across populations with features that were never observed and those with different overall sample missingness. There was a significantly higher percentage of positive (mortality) outcomes in patients with $\leq$25of features never observed across all three datasets, where the prevalence of outcomes increased more than double compared to those with higher feature-wise missingness. Similar trends were also observed in sample-wise missingness, suggesting that the missingness pattern is personalized and could inform the underlying patient’s health status for various personalized medical applications. Imputation methods that consider individual missingness patterns, such as IGNITE, could lead to more accurate predictions and tailored treatments. Moreover, our feature-wise missingness experiments showcased IGNITE’s power to generate features that were never observed in patients. In traditional statistical approaches, many patients with such never-observed features are generally omitted/removed, leading to a waste of data or a reduction in model performance. However, with IGNITE, such features can be generated based on other readily available information about an individual. We believe that a flexible generative approach proposed by IGNITE can open doors for a new way of rethinking missingness imputation as a generative task, making various new applications utilizing generative models.
@@
Another strength of our proposed work is that IGNITE is designed to generate imputations conditioned on individualized patient observed and missingness patterns and dynamics, making the generated imputations of high quality and more aligned with personalized medicine applications. Generating digital twins to drive personalized medicine applications requires understanding and representing patient data beyond observed values, which IGNITE incorporates into its generative process. Lastly, IGNITE was trained and tested across three large-scale EHRs, outperforming state-of-the-art imputation models across a series of experiments. Other experiments, such as the reconstruction experiments, demonstrate IGNITE’s robustness and ability to recover original patient values even when random missingness is introduced to patient records. This robust performance across datasets and experimental settings showcases IGNITE’s robustness in learning individualized representations and generating high-quality imputations.
Overall, this study has some limitations. IGNITE and the other imputation models were tested only on retrospective ICU datasets. In future work, we plan to investigate IGNITE’s ability to learn individualized patient representations with higher sparsity and missingness patterns in primary care and wearable data. Furthermore, while our work expands on the evaluation metrics proposed in the literature, IGNITE was only evaluated in prediction tasks involving mortality as limited by the available datasets such as Physionet 2012. In future work, we aim to investigate the impact of imputations on various tasks in healthcare settings, such as treatment recommendation and phenotyping. Another area of future work involves the theoretical analysis of the various missingness patterns in the data to better understand the strengths and limitations of the plethora of time-series techniques. In addition, the datasets selected for the evaluation of the IGNITE model were chosen based on their open-access nature to ensure that our results are transparent and reproducible. These ICU datasets are commonly used in healthcare analytics research, allowing us to benchmark our results against well-established methods in the field. Although the current study focuses on ICU datasets due to the availability and richness of their clinical data, we recognize that this may limit the perceived applicability of our method. However, it is important to note that the principles underlying the IGNITE model are broadly applicable to other types of clinical data. Future research will aim to apply IGNITE to additional datasets from diverse medical settings, demonstrating its broader applicability and exploring its potential in other areas of clinical practice. This expansion will also allow us to test the model’s robustness across different healthcare environments, addressing any unique challenges they present, and supporting the reproducibility of our findings. Another potential area of future work is to include more benchmarks that might be relevant in medical statistics such as Zhang et al., (2020); Sun, (2019). For the sake of this work, we selected traditional statistical methods such as MICE and LOCF alongside advanced deep learning models, to provide a comprehensive perspective on imputation effectiveness. This selection strategy ensures a robust comparison across different methodological approaches, demonstrating the advances deep learning offers over traditional benchmarks, particularly when learning individualized imputations. Lastly, while deep learning models like IGNITE require more computational resources during training compared to traditional methods, they significantly expedite the inference process. In our analysis, the improvements in data quality and model performance using IGNITE are statistically significant, validating the trade-off between increased training time and enhanced operational efficiency in practical applications.
@@ 1.4 Notations
We denote a multivariate time-series EHRs dataset as $\mathbf{D}=\left\{\left(\mathbf{x}_{i,1:T_{i}}\right)\right\}_{i=1}^{N}$, which includes a set of individual patient records indexed by $i\in\{1,2,\ldots N\}$. Each patient record comprises two main components: time-series and static components. The time-series component includes continuous time-series of physiological data and the discrete time-series of treatment data, recorded over time steps $\mathbf{T}=(\mathbf{t}_{1},\ldots,{t}_{N})$. Each record has a corresponding static outcome label $\mathbf{y}=(\mathbf{y}_{1},\ldots,{y}_{N})$ and static demographic information, indicating the age and sex of the patient. For each individual patient time-series matrix, we extract a binary mask indicating the missingness of each value, where 0 and 1 denote the missing value and the observed value, respectively.
@@ 1.5.1 Dual Variational-AutoEncoders with Dual-Stage Attention
We utilize a pair of VAEs, each with an encoder and a decoder, to map the multi-variate time-series into reversible low-dimensional dense representations with no missingness. We use long short-term memory (LSTM) Hochreiter and Schmidhuber, (1997) neural networks for the architecture of VAE. The first VAE is trained to reconstruct the multivariate time-series by calculating the evidence lower bound (ELBO) on the observed values only and imputing the missing values with zero, an approach adopted by (Nazabal et al.,, 2020; Fortuin et al.,, 2020). The second VAE, on the other hand, generates the data based on the ELBO calculated on the full data that is first imputed with the last value carried forward and masked with the Individualized Missingness Mask (IMM) that is discussed in section 1.5.2. We refer to the first and second VAE as $VAE^{\mathcal{OO}}$and $VAE^{\mathcal{IMM}}$, respectively. Each LSTM encoder is complemented with input attention highlighting feature importance. Specifically, the attention component computes the weights for each feature conditioned on the encoder’s hidden state of the previous time step. The computed weights are obtained using a deterministic fully connected network, by referring to the previous hidden state $\mathbf{h}_{t-1}$ and the cell state $\mathbf{s}_{t-1}$ in the encoder LSTM unit with:
@@
where $\alpha_{t}^{k}$ is the attention weight of feature $k$ at time $t$. The learnable parameters are denoted in the form of $\mathbf{v}_{e}\in\mathbb{R}^{T},\mathbf{W}_{e}\in\mathbb{R}^{T\times 2m}$ and $\mathbf{U}_{e}\in\mathbb{R}^{T\times T}$. The learned attention weights are passed to a softmax function, followed by a dense network trained with the encoder parameters. To this end, the learned weights are multiplied by the input features as shown in
@@
The LSTM encoder unit are fed with $\tilde{\mathbf{x}}_{t}$ instead of ${\mathbf{x}}_{t}$, giving more attention to important features to encode a better representation of the time-series input. After encoding the data with feature-wise attention, a temporal attention mechanism is used in the decoder to select relevant encoder hidden states across all time steps. To do so, an attention weight is calculated for each encoder hidden state at time $t$. The calculation of the attention weight is based on the previous hidden state at time $t-1$. Specifically, the attention weight of each encoder hidden state at time $t$ is calculated based on the previous hidden state of the decoder $\mathbf{d}_{t-1}\in\mathbb{R}^{p}$ and the cell state of the LSTM unit $\mathbf{s}_{t-1}^{\prime}\in\mathbb{R}^{p}$ with
@@
where $\left[\mathbf{d}_{t-1};\mathbf{s}_{t-1}^{\prime}\right]\in\mathbb{R}^{2p}$ is a concatenation of the previous hidden state and the cell state of the LSTM unit. $\mathbf{v}_{d}\in\mathbb{R}^{m}$, $\mathbf{W}_{d}\in\mathbb{R}^{m\times 2p}$ and $\mathbf{U}_{d}\in\mathbb{R}^{m\times m}$ are parameters to learn. The attention weight $\beta_{t}^{i}$ represents the importance of the $i$-th encoder hidden state at time $t$. Since each encoder hidden state $\mathbf{h}_{i}$ is mapped to a temporal component of the input, the attention mechanism computes the context vector $\mathbf{c}_{t}$ as a weighted sum of all the encoder hidden states $\left\{\mathbf{h}_{1},\mathbf{h}_{2},\cdots,\mathbf{h}_{T}\right\}$,
@@
We note that the context vector $\mathbf{c}_{t}$ is distinct at each time step. Once we get the weighted summed context vectors. The decoded data are based on the stacked weighted content vectors, allowing for better representation over time.